Romeo is throwing pebbles gently up to Juliet's wi
Answer #1

final vy = 0 = vy - 9.81 t

vy = 9.81 t
final y = 9.2m = vyt - g/2 t^2= vy t - 4.905 t^2

9.20 m = 9.81t* t - 4.905 t^2 = 4.905 t^2
t = sqrt(9.20m/4.905m/s^2) = 1.369 s
final vx = original vx =vx
final x = 9.4 m = vx t
at window
vx = 9.4m /t =9.4/1.369 = 6.866 m/s ..................ans
vy = 0

so, the pebble is going to hit with 6.866 m/s speed